# A Comprehensive Survey on Underwater Image Enhancement Based on Deep Learning

### v3

@@ Abstract
Index Terms:
Underwater Image Enhancement, Quality Degradation, Color Distortion, Light Attenuation.
@@ I Introduction
Solutions have been proposed. On the one hand, we summarize the physical models, data generation methods, and evaluation metrics which are constructed to serve as the basis for research. On the other hand, numerous network structures and training strategies are classified and discussed. In order to know which methods have been widely explored, we need to understand the similarities and differences between them. Within our perspective, existing research can be subjectively viewed as being carried out from six different aspects, namely network architecture, learning strategy, learning stage, assistance task, domain perspective and disentanglement &\&& fusion. To facilitate future research that can easily build on existing work, we provide a comprehensive review of existing algorithms. According to the main contributions of different algorithms, we divide the existing work into 6 first-level categories. Then, for each first-level category, we give the corresponding second-level category. The taxonomy of algorithms is shown in Table I.
@@ I-B A guide for reading this survey
Disentanglement &\&& Fusion Physical Embedding
@@ III UIE Methods
According to main contributions of each paper, we divide UIE algorithms into six categories, namely (i) Section III-A Network Architecture, (ii) Section III-B Learning Strategy, (iii) Section III-C Learning Stage, (iv) Section III-D Assistance Task, (v) Section III-E Domain Perspective and (vi) Section III-F Disentanglement &\&& Fusion, respectively. Details are as follows.
@@
  • Disentanglement &\&& Fusion. Considering the degradation properties of underwater images, disentanglement and fusion are used as two effective ways to improve the interpretability of the model. Physical models and Retinex models are usually used as the theoretical basis for disentanglement. Color space fusion, water type fusion and multi-input fusion are common solutions for fusion.
@@ III-B2 Rank Learning
$\mathcal{L}(s_{n},s_{m})=\begin{cases}\max{(0,(s_{m}-s_{n})+\epsilon)},q_{n}>q_{m},\\ \max{(0,(s_{n}-s_{m})+\epsilon)},q_{n}<q_{m},\\ \end{cases}$ (11)
@@ III-E1 Knowledge Transfer
$\begin{split}\mathcal{L}^{img}&=\mathbb{E}_{x_{st}}[D^{img}(x_{st})]-\mathbb{E}_{x_{r}}[D^{img}(x_{r})]\\ &=+\lambda_{img}\mathbb{E}_{\hat{I}}(||\nabla_{\hat{I}}D^{img}(\hat{I})||_{2}-1)^{2},\end{split}$ (19)
$\begin{split}\mathcal{L}^{out}&=\mathbb{E}_{y_{st}}[D^{out}(y_{st})]-\mathbb{E}_{y_{r}}[D^{out}(y_{r})]\\ &+\lambda_{out}\mathbb{E}_{\hat{I}}(||\nabla_{\hat{I}}D^{out}(\hat{I})||_{2}-1)^{2},\end{split}$ (20)
@@
$\begin{split}\mathcal{L}^{feat}&=\mathbb{E}_{x_{st}}[D^{feat}(\mathcal{G}_{r}(x_{st}))]-\mathbb{E}_{x_{r}}[D^{feat}(\mathcal{G}_{r}(x_{r}))]\\ &+\lambda_{feat}\mathbb{E}_{\hat{I}}(||\nabla_{\hat{I}}D^{feat}(\hat{I})||_{2}-1)^{2},\end{split}$ (21)
@@ III-E2 Domain Translation
$\begin{cases}\mathcal{L}^{img}=||x-\dot{x}||_{1}+||y-\dot{y}||_{1},\\ \mathcal{L}^{sty}=||s_{x}-\dot{s_{x}}||_{1}+||s_{y}-\dot{s_{y}}||_{1},\\ \mathcal{L}^{con}=||c_{x}-\dot{c_{x}}||_{1}+||c_{y}-\dot{c_{y}}||_{1},\\ \end{cases}$ (22)
@@ III-F4 Water Type Focus
$\begin{cases}\mu=\frac{1}{HW}x_{n},\\ \Sigma=\frac{1}{HW}(x_{n}-\mu)(x_{n}-\mu)^{T}+\alpha{I},\\ \end{cases}$ (30)
@@ IV-B Comparison among DL-UIE Algorithms
Evaluation of Fitting Ability on Full-Reference Benchmark Datasets. Table LABEL:tab:full_reference_results shows the results obtained by various DL-UIE algorithms on datasets with paired data. Overall, the quantitative evaluation metrics obtained by UIE-DM are the best. However, on most datasets, there is no significant difference in the quantitative evaluation results (i.e., PSNR and SSIM) obtained by the top 5 algorithms. What is particularly noteworthy is that for EUVP-D, EUVP-I and EUVP-S, the PSNR and SSIM of the top 10 algorithms are approaching the same level. The visualization results in Fig. 8 also illustrate that multiple DL-UIE algorithms achieve similar visual effects. Not only that, the color histogram also shows that the results obtained by different algorithms are close in terms of pixel statistics. For these UIE algorithms, the representation capabilities of the neural networks designed with different architectures may be very close. This phenomenon implies that the pursuit of higher PSNR and SSIM may be difficult.
Evaluation of Generalization Ability on No-Reference Benchmark Datasets. We tested two kinds no-reference metrics. The first is the manually calculated metrics UIQM and UCIQE, which are shown in Table LABEL:tab:non_reference_results. The UIQM and UCIQE values of different DL-UIE algorithms given in Table LABEL:tab:non_reference_results show that UGAN achieves the overall best performance. Meanwhile, in terms of the evaluation of generalization ability, the UIQM and UCIQE values obtained by different algorithms are close. The second is URANKER obtained by data-driven training, which is given in Fig. 9. The best URANKER values are obtained by M (UIE-DM), K (CECF), E (UGAN), and I (STSC), respectively. The second best URANKER values are obtained by K (CECF), I (STSC), H (SCNet), M (UIE-DM) respectively. For the evaluation of URANKER, UIE-DM has the best overall evaluation effect. Considering Table LABEL:tab:non_reference_results and Fig. 9 simultaneously, the generalization ability of UGAN is the best. Fig. 10 shows the visual enhancement results of the evaluation of generalization ability. There are obvious differences in the visual results and pixel histograms of images obtained by different algorithms. Some enhanced images still show blue-green of green effect.

### v4

@@ I Introduction
Solutions have been proposed. On the one hand, we summarize the physical models, data generation methods, and evaluation metrics which are constructed to serve as the basis for research. On the other hand, numerous network structures and training strategies are classified and discussed. In order to know which methods have been widely explored, we need to understand the similarities and differences between them. Within our perspective, existing research can be subjectively viewed as being carried out from six different aspects, namely network architecture, learning strategy, learning stage, assistance task, domain perspective and disentanglement &\& fusion. To facilitate future research that can easily build on existing work, we provide a comprehensive review of existing algorithms. According to the main contributions of different algorithms, we divide the existing work into 6 first-level categories. Then, for each first-level category, we give the corresponding second-level category. The taxonomy of algorithms is shown in Table I.
@@ I-B A guide for reading this survey
Disentanglement &\& Fusion Physical Embedding
@@ III UIE Methods
According to main contributions of each paper, we divide UIE algorithms into six categories, namely (i) Section III-A Network Architecture, (ii) Section III-B Learning Strategy, (iii) Section III-C Learning Stage, (iv) Section III-D Assistance Task, (v) Section III-E Domain Perspective and (vi) Section III-F Disentanglement &\& Fusion, respectively. Details are as follows.
@@
  • Disentanglement &\& Fusion. Considering the degradation properties of underwater images, disentanglement and fusion are used as two effective ways to improve the interpretability of the model. Physical models and Retinex models are usually used as the theoretical basis for disentanglement. Color space fusion, water type fusion and multi-input fusion are common solutions for fusion.
@@ III-B2 Rank Learning
[…]
@@ III-E1 Knowledge Transfer
[…]
@@
[…]
@@ III-E2 Domain Translation
[…]
@@ III-E3 Diversified Output
TABLE II: Quantitative results with supervised training on full-reference benchmarks in terms of PSNR and SSIM. Best results are in red and second best results are in blue.
Method UIEB LSUI EUVP-D EUVP-I EUVP-S UFO-120
$uparrow PSNR\uparrow SSIM\uparrow$ $uparrow PSNR\uparrow SSIM\uparrow$ $uparrow PSNR\uparrow SSIM\uparrow$
UWNet [99] 17.36 0.7948 22.01 0.8489 21.36 0.8887 22.54 0.8115 25.45 0.8616 25.02 0.8505
PhysicalNN [15] 18.08 0.7693 22.52 0.8468 21.37 0.8851 23.68 0.8108 26.07 0.8507 25.68 0.8414
FUnIEGAN [53] 19.12 0.8321 24.23 0.8579 21.04 0.8471 23.06 0.7731 27.55 0.8695 26.79 0.8378
WaterNet [74] 21.04 0.8601 22.74 0.8560 21.40 0.8881 23.99 0.8325 23.43 0.8287 23.84 0.8188
ADMNNet [141] 20.77 0.8709 26.11 0.9051 22.29 0.9063 24.65 0.8424 27.11 0.8893 26.67 0.8757
UColor [71] 20.13 0.8769 24.03 0.8855 22.43 0.9088 24.43 0.8462 27.65 0.9036 26.95 0.8877
UGAN [28] 21.37 0.8534 26.93 0.8983 22.02 0.8850 24.71 0.8276 28.26 0.8851 27.61 0.8672
SGUIE [105] 21.12 0.8882 27.77 0.9129 22.06 0.8927 24.68 0.8356 28.83 0.8934 28.30 0.8784
UIE-WD [94] 20.10 0.8617 26.81 0.9088 22.21 0.9080 24.56 0.8365 28.43 0.8990 27.48 0.8756
SCNet [30] 21.37 0.8981 27.47 0.9111 22.40 0.9028 24.69 0.8267 28.29 0.8850 27.34 0.8690
STSC [122] 21.05 0.8661 27.67 0.9108 22.45 0.9084 25.27 0.8484 29.05 0.8956 28.16 0.8847
U-Trans [102] 20.39 0.8034 27.48 0.9013 22.15 0.8941 24.56 0.8309 28.46 0.8843 27.60 0.8525
CECF [19] 21.82 0.8866 27.27 0.9124 22.21 0.9032 24.86 0.8567 27.13 0.8879 26.75 0.8729
Semi-UIR [47] 22.79 0.9088 25.44 0.8801 22.31 0.9059 24.46 0.8170 26.07 0.8402 28.41 0.8837
UIE-DM [117] 23.03 0.9099 29.56 0.9194 22.13 0.9066 25.14 0.8499 26.12 0.8757 26.10 0.8464
HCLR-Net [168] 22.09 0.8976 27.94 0.9169 22.43 0.9075 24.92 0.8314 29.23 0.8928 28.85 0.8818
GUPDM [98] 21.76 0.8697 27.34 0.9116 22.22 0.8739 24.91 0.8330 28.26 0.8904 27.81 0.8828
@@
TABLE III: Quantitative results on no-reference benchmarks in terms of UIQM and UCIQE. Best results are in red and second best results are in blue.
Method Challenge-60 U45 UCCS EUVP-330
$uparrow UIQM\uparrow UCIQE\uparrow$ $uparrow UIQM\uparrow UCIQE\uparrow$
Distortion 3.041 0.5187 2.347 0.5306 1.698 0.4949 2.264 0.5268
UWNet 3.832 0.5061 3.792 0.5183 3.194 0.4821 3.429 0.5160
PhysicalNN 3.946 0.5444 3.797 0.5578 3.317 0.5067 3.368 0.5475
FUnIEGAN 4.325 0.5819 4.372 0.6174 4.086 0.5481 3.636 0.5964
WaterNet 4.195 0.5850 4.167 0.5960 4.158 0.5654 3.810 0.6026
ADMNNet 4.190 0.5649 4.277 0.5836 4.062 0.5477 3.748 0.5775
UColor 3.992 0.5565 4.144 0.5793 3.583 0.5357 3.552 0.5694
UGAN 4.480 0.5891 4.501 0.6014 4.391 0.5654 3.899 0.6153
SGUIE 4.221 0.5791 4.256 0.6000 3.967 0.5525 3.663 0.5927
UIE-WD 4.102 0.5654 4.164 0.6004 3.711 0.5264 3.557 0.5810
SCNet 4.259 0.5813 4.456 0.6005 4.386 0.5727 3.857 0.5991
STSC 4.261 0.5751 4.343 0.5988 4.004 0.5516 3.717 0.5923
U-Trans 4.063 0.5604 4.120 0.5914 3.819 0.5365 3.525 0.5746
CECF 4.127 0.5728 4.354 0.5918 4.286 0.5531 3.681 0.5903
Semi-UIR 4.098 0.5761 4.431 0.5999 4.049 0.5499 3.660 0.5974
UIE-DM 3.874 0.5639 4.097 0.5974 4.166 0.5673 3.652 0.5984
HCLR-Net 4.100 0.5656 4.293 0.6027 3.935 0.5376 3.680 0.5861
GUPDM 4.276 0.5652 4.399 0.5897 4.108 0.5300 4.017 0.5823
@@ III-F4 Water Type Focus
[…]
@@ IV-B Comparison among DL-UIE Algorithms
Evaluation of Fitting Ability on Full-Reference Benchmark Datasets. Table II shows the results obtained by various DL-UIE algorithms on datasets with paired data. Overall, the quantitative evaluation metrics obtained by UIE-DM are the best. However, on most datasets, there is no significant difference in the quantitative evaluation results (i.e., PSNR and SSIM) obtained by the top 5 algorithms. What is particularly noteworthy is that for EUVP-D, EUVP-I and EUVP-S, the PSNR and SSIM of the top 10 algorithms are approaching the same level. The visualization results in Fig. 8 also illustrate that multiple DL-UIE algorithms achieve similar visual effects. Not only that, the color histogram also shows that the results obtained by different algorithms are close in terms of pixel statistics. For these UIE algorithms, the representation capabilities of the neural networks designed with different architectures may be very close. This phenomenon implies that the pursuit of higher PSNR and SSIM may be difficult.
Evaluation of Generalization Ability on No-Reference Benchmark Datasets. We tested two kinds no-reference metrics. The first is the manually calculated metrics UIQM and UCIQE, which are shown in Table III. The UIQM and UCIQE values of different DL-UIE algorithms given in Table III show that UGAN achieves the overall best performance. Meanwhile, in terms of the evaluation of generalization ability, the UIQM and UCIQE values obtained by different algorithms are close. The second is URANKER obtained by data-driven training, which is given in Fig. 9. The best URANKER values are obtained by M (UIE-DM), K (CECF), E (UGAN), and I (STSC), respectively. The second best URANKER values are obtained by K (CECF), I (STSC), H (SCNet), M (UIE-DM) respectively. For the evaluation of URANKER, UIE-DM has the best overall evaluation effect. Considering Table III and Fig. 9 simultaneously, the generalization ability of UGAN is the best. Fig. 10 shows the visual enhancement results of the evaluation of generalization ability. There are obvious differences in the visual results and pixel histograms of images obtained by different algorithms. Some enhanced images still show blue-green of green effect.